# Symphony-MoE: Harmonizing Disparate Pre-trained Models into a Coherent Mixture-of-Experts

### v1

@@ Abstract
Mixture-of-Experts (MoE) models enable scalable performance by activating large parameter sets sparsely, minimizing computational overhead. To circumvent the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts. However, this limits expert diversity, as all experts originate from a single pre-trained dense model. This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Llama2-Chat and Code Llama). A key challenge lies in the fact that these source models occupy disparate, dissonant regions of the parameter space, making direct upcycling prone to severe performance degradation. To overcome this, we propose Symphony-MoE, a novel two-stage framework designed to harmonize these models into a single, coherent expert mixture. First, we establish this harmony in a training-free manner: we construct a shared backbone via a layer-aware fusion strategy and, crucially, alleviate parameter misalignment among experts using activation-based functional alignment. Subsequently, a single lightweight stage of router training coordinates the entire architecture. Experiments demonstrate that our method successfully integrates experts from heterogeneous sources, achieving an MoE model that significantly surpasses baselines in multi-domain tasks and out-of-distribution generalization.
@@ 1 Introduction
However, existing upcycling approaches such as “Branch-Train-Mix” (Sukhbaatar et al. 2024) share a fundamental limitation: reliance on a single pre-trained checkpoint (see Figure 1). Consequently, all experts originate from identical parameter space, regardless of subsequent specialization. This “single-origin” constraint intrinsically caps expert diversity. In practice, a large number of high-quality, identically architected models—specialized through distinct pre-training objectives, data corpora, or versions—exist, such as the dialogue-optimized Llama2-Chat (Touvron et al. 2023) and the code-centric Code Llama (Roziere et al. 2023). These models capture unique knowledge and capabilities within their domains, occupying valuable yet heterogeneous regions of the parameter space. Directly leveraging such specialized models as distinct experts offers a promising avenue for constructing more powerful and versatile MoEs, but remains infeasible under current frameworks.
Figure 2: Overview of the Symphony-MoE construction pipeline. In Stage 1, disparate source models are harmonized without training: non-FFN layers are merged into a shared backbone using techniques such as SLERP, while FFN layers are aligned through activation-based neuron permutation. In Stage 2, only the router is trained while keeping the backbone and expert parameters frozen, enabling coordination among the now-compatible experts.
@@
The second stage implements lightweight global coordination. We freeze the expert layers and shared backbone to train the router independently. To prevent expert collapse and ensure balanced utilization, this stage critically employs an auxiliary loss function promoting load balancing across experts.
@@
(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified Mixture-of-Experts (MoE) system. We identify parameter space misalignment as a key barrier to this process..
(2) We introduce a lightweight framework that addresses expert incompatibility without gradient-based training, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of upcycling disparate pre-trained models into unified expert systems.
@@ 2.2 MoE Construction via Upcycling
To mitigate the immense cost of training MoEs from scratch, the upcycling paradigm has emerged as a practical alternative. The vanilla approach, or Sparse Upcycling (Komatsuzaki et al. 2022; He et al. 2024), involves converting a pre-trained dense model into an MoE by replicating its feed-forward network (FFN) layers to form multiple, initially identical experts. While this provides a strong initialization, it suffers from a knowledge bottleneck and a lack of initial expert diversity, as all experts originate from a single source.
@@ 2.3 Model Merging
The challenge of combining disparate pre-trained models is the central focus of the model merging field (Tao et al. 2024). To merge truly “stranger” models, advanced techniques explicitly resolve misalignment. Model Soups (Wortsman et al. 2022) averages the weights of models fine-tuned from the same checkpoint, assuming they remain in the same well-connected loss basin. Methods like PLeaS (Ito, Yamada, and Kumagai 2024) and Git Re-Basin (Ainsworth, Hayase, and Srinivasa 2022) find an optimal permutation matrix to align neurons layer-wise before averaging. TIES-Merging (Yadav et al. 2023) resolves interference between task vectors by trimming redundant updates and electing a dominant sign for parameter changes. These methods, however, are general-purpose algorithms designed to produce a single, dense consensus model. They are not tailored for the specific architectural needs of MoE construction, such as preserving expert diversity or informing the initialization of the router.
Our work bridges this gap. We integrate permutation-based alignment directly into the MoE construction pipeline, repurposing these techniques for a new architectural goal. We use alignment to make disparate experts functionally compatible. This ensures they can operate within a shared coordinate system while preserving their individual specializations.
@@ 3 Methods
We introduce a two-stage framework to effectively assemble a MoE model from disparate, pre-trained source models. The first stage operates entirely without training, resolving the critical parameter space misalignment problem through training-free functional alignment. The second stage then trains a router that coordinates the now-compatible experts, unlocking their collective capabilities.
@@ 3.1 Preliminaries and Problem Formulation
Our work addresses the novel challenge of constructing an MoE model from multiple, disparate pre-trained source models. Formally, we are given a set of $N$ source models, $\{M_{1},M_{2},\ldots,M_{N}\}$. These models share an identical architecture but possess distinct sets of parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ as a result of being trained on different datasets $\{D_{1},D_{2},\ldots,D_{N}\}$ , for different tasks, or with different optimization objectives. The goal is to construct a single, powerful MoE model, $M_{MoE}$, by leveraging the FFN layers of the source models as its experts. The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible numerical and semantic spaces.
@@ 3.2 Stage 1: Training-Free Functional Alignment
Our first stage assembles a cohesive MoE model from diverse parent models entirely without training. This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN layers, which will serve as the experts.
@@
1. Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling 5% from each instruction fine-tuning dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training. Each instruction in $D_{\mathrm{cal}}$ is passed through both the anchor model $M_{1}$ and a target model $M_{i}$ to extract their FFN activation matrices at layer $l$, denoted as $A_{1}^{(l)}$ and $A_{i}^{(l)}$, respectively.
@@
3. Weight Remapping: We apply the computed permutation $P_{i}^{(l)}$ to the FFN weights (Nasery et al. 2025). This involves permuting the output dimension of the first linear layer’s weights $W_{up}$ and the input dimension of the second’s $W_{down}$ to maintain functional equivalence:
$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (3)
@@
We analyzed the computational cost to verify the scalability. The analysis can be found in the Appendix.
3.3 Stage 2: Activating the MoE via Lightweight Router Training
In Stage 2, we integrate the constructed shared backbone with the expert models and introduce a randomly initialized router, implemented as a simple linear layer. The expert layers and shared backbone—which encode knowledge from the source models—are frozen, and only the router is trained. The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is reused for this purpose, enabling the router to learn how to coordinate the experts.
We adopt a top-2 routing mechanism and train the router using a composite loss function:
@@ 3.3 Stage 2: Activating the MoE via Lightweight Router Training
where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term. The coefficient $\lambda$ is fixed at 0.01 in all experiments. $\mathcal{L}_{bal}$ encourages uniform expert utilization, which mitigates expert collapse and allows the model to fully leverage the diversity of its specialists.
Notably, the router is not provided with explicit supervision indicating which expert to select for each token, as there is no ground-truth mapping between tokens and experts. Instead, it learns expert assignment implicitly through training on instruction–response pairs in $D_{\mathrm{cal}}$.
@@ 4.1.1 Build Disparately Initialized Dense Models
MMLU GSM8K BBH HumanEval TruthfulQA Avg. MedCQA
@@
Table 1: Performance comparison of dense (Qwen2/2.5 0.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
@@
Table 2: Performance comparison of dense (Qwen2/2.5 1.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model.
@@ 4.1.2 Evaluation Datasets and Metrics
We evaluate our method on five in-distribution datasets: MMLU (General)  (Hendrycks et al. 2021), GSM8K (Math) (Cobbe et al. 2021), BBH (General, Reasoning)  (Suzgun et al. 2022), HumanEval (Code) (Chen et al. 2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022). We also include an out-of-distribution benchmark, MedCQA (Medicine) (Shoham and Rappoport 2024).
@@ 4.2 Implementation Details
Our experiments were conducted on 8 NVIDIA V100 GPUs. The dense model was implemented using LLaMA-Factory (Zheng et al. 2024), and trained for for 2 epoches. Detailed training configurations are provided in the Appendix.
In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.404 million tokens. To ensure consistent alignment and integration across expert models, we fixed a general-purpose expert model $M_{1}$ as the anchor model. To ensure that the knowledge of each expert is equally integrated and to avoid introducing any bias, the weights for merging the 4 dense models are set to 0.25.
In Stage 2: Router Training, we set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2. The learning rate is fixed at 5e-5, and optimization is performed using AdamW (Zhuang et al. 2022) with a maximum gradient norm of 1.0.
@@ 4.4 Analysis of Inter-Expert Functional Specialization
To quantitatively assess the effectiveness of our functional alignment method, we compute the pairwise functional specialization between expert layers using Centered Kernel Alignment (CKA) (Cortes, Mohri, and Rostamizadeh 2012), a widely adopted technique for comparing neural network representations. A lower CKA score indicates greater functional specialization. We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; and (c) our proposed method employing activation-based functional alignment.
As shown in Figure 3, the naive merging strategy leads to a severe representational collapse. Experts trained independently develop distinct internal coordinate systems; merging them without alignment forces the model to conflate functionally unrelated neurons. This results in inflated inter-expert CKA scores (e.g., 0.65–0.75), signaling that unique functional fingerprints of the experts have been blurred into a redundant subspace. In such a scenario, the router lacks the discriminative basis to perform meaningful, specialized dispatching.
In contrast, our method explicitly resolves this misalignment by permuting neurons to form a functionally coherent coordinate system prior to merging. Figure 3(c) demonstrates that this alignment dramatically restores inter-expert specialization, with CKA scores returning to near-optimal levels comparable to those of the original experts (Figure 3(a)). This provides empirical evidence that our alignment strategy effectively mitigates parameter space misalignment.
@@ 4.7 Ablation Analysis
The results highlight the critical role of functional alignment: removing it causes catastrophic performance collapse, confirming that resolving parameter space inconsistencies is fundamental to the framework. Replacing SLERP with linear averaging leads to substantial degradation, supporting our hypothesis that the geometry of attention weights is essential for preserving functionality. While simplifying the embedding fusion yields only a moderate drop, it nonetheless indicates the value of layer-aware merging in the shared backbone. Notably, using biased calibration data results in the most severe performance decline after removing alignment, underscoring the importance of domain diversity in constructing $D_{\mathrm{cal}}$ for effective inter-expert coordination.

### v2

@@ Abstract
Mixture-of-Experts (MoE) models enable scalable performance by activating large parameter sets sparsely, minimizing computational overhead. To mitigate the prohibitive cost of training MoEs from scratch, recent work employs upcycling, reusing a single pre-trained dense model by replicating its feed-forward network (FFN) layers into experts. However, this limits expert diversity, as all experts originate from a single pre-trained dense model. This paper addresses this limitation by constructing powerful MoE models using experts sourced from multiple identically-architected but disparate pre-trained models (e.g., Qwen2.5-Coder and Qwen2). A key challenge lies in the fact that these source models occupy disparate, dissonant regions of the parameter space, making direct upcycling prone to severe performance degradation. To overcome this, we propose Symphony-MoE, a novel two-stage framework designed to harmonize these models into a single, coherent expert mixture. First, we establish this harmony in a training-free manner: we construct a shared backbone via a layer-aware fusion strategy and, crucially, alleviate parameter misalignment among experts using activation-based functional alignment. Subsequently, a stage of post-training coordinates the entire architecture. Experiments demonstrate that our method successfully integrates experts from heterogeneous sources, achieving an MoE model that significantly surpasses baselines in multi-domain tasks and out-of-distribution generalization.
@@ 1 Introduction
However, existing upcycling approaches such as “Branch-Train-Mix” (Sukhbaatar et al. 2024) share a fundamental limitation: reliance on a single pre-trained checkpoint (see Figure 1). Consequently, all experts originate from an identical parameter space, regardless of subsequent specialization. This “single-origin” constraint intrinsically caps expert diversity. In practice, a large number of high-quality, identically architected models—specialized through distinct pre-training objectives, data corpora, or versions—exist, such as the dialogue-optimized Llama2-Chat (Touvron et al. 2023) and the code-centric Code Llama (Roziere et al. 2023). These models capture unique knowledge and capabilities within their domains, occupying valuable yet heterogeneous regions of the parameter space. Directly leveraging such specialized models as distinct experts offers a promising avenue for constructing more powerful and versatile MoEs, but remains infeasible under current frameworks.
Figure 2: Overview of the Symphony-MoE construction pipeline. In Stage 1, disparate source models are harmonized without training: non-FFN layers are merged into a shared backbone using techniques such as SLERP, while FFN layers are aligned through activation-based neuron permutation. In Stage 2, the experts, router, and shared backbone undergo post-training to enable coordination among the now-compatible components.
@@
The second stage performs global coordination through post-training. To prevent expert collapse and encourage balanced expert utilization, this stage introduces a load-balancing loss that promotes uniform routing across experts.
@@
(1) We investigate the problem of upcycling multiple identically architected yet independently pretrained models into a unified MoE model. We identify parameter space misalignment as a key barrier to this process.
(2) We introduce a harmonization framework that addresses expert incompatibility, combining layer-aware fusion and activation-driven functional alignment to preserve the specialized knowledge embedded in each source model.
(3) Through comprehensive experiments on in-distribution and out-of-distribution datasets, we show that our method achieves strong performance and generalizes well to out-of-distribution settings, demonstrating the feasibility of harmonizing disparate pre-trained models into a coherent MoE.
@@ 2.2 MoE Construction via Upcycling
To mitigate the immense cost of training MoEs from scratch, the upcycling paradigm has emerged as a practical alternative. The vanilla approach, or Sparse Upcycling (Komatsuzaki et al. 2022; He et al. 2024; Zhao et al. 2024), involves converting a pre-trained dense model into an MoE by replicating its feed-forward network (FFN) layers to form multiple, initially identical experts. While this provides a strong initialization, it suffers from a knowledge bottleneck and a lack of initial expert diversity, as all experts originate from a single source.
@@ 2.3 Model Merging
The challenge of combining disparate pre-trained models is the central focus of the model merging field (Tao et al. 2024; Imfeld et al. 2023). To merge truly “stranger” models, advanced techniques explicitly resolve misalignment. Model Soups (Wortsman et al. 2022) averages the weights of models fine-tuned from the same checkpoint, assuming they remain in the same well-connected loss basin. Methods like PLeaS (Ito, Yamada, and Kumagai 2024) and Git Re-Basin (Ainsworth, Hayase, and Srinivasa 2022) find an optimal permutation matrix to align neurons layer-wise before averaging. TIES-Merging (Yadav et al. 2023) resolves interference between task vectors by trimming redundant updates and electing a dominant sign for parameter changes. These methods, however, are general-purpose algorithms designed to produce a single, dense consensus model. They are not tailored for the specific architectural needs of MoE construction, such as preserving expert diversity or informing the initialization of the router.
Our work bridges this gap. We integrate permutation-based alignment into the MoE construction pipeline, repurposing these techniques for a new goal. We use alignment to make disparate experts functionally compatible. This ensures they can operate within a shared coordinate system while preserving their individual specializations.
@@ 3 Methods
We introduce a two-stage framework to effectively assemble a MoE model from disparate, pre-trained source models. The first stage operates entirely without training, resolving the critical parameter space misalignment problem through training-free functional alignment. The second stage then coordinates the now-compatible experts, unlocking their collective capabilities.
@@ 3.1 Preliminaries and Problem Formulation
Our work addresses the novel challenge of constructing an MoE model from multiple, disparate pre-trained source models. Formally, we are given a set of $N$ source models, $\{M_{1},M_{2},\ldots,M_{N}\}$. These models share an identical architecture but possess distinct sets of parameters $\{\theta_{1},\theta_{2},\ldots,\theta_{N}\}$ as a result of being trained on different datasets $\{D_{1},D_{2},\ldots,D_{N}\}$, for different tasks, or with different optimization objectives. The goal is to construct a single, powerful MoE model, $M_{MoE}$, by leveraging the FFN layers of the source models as its experts. The fundamental barrier to this goal is the severe parameter space misalignment between the models, where the parameters $\theta_{i}$ and $\theta_{j}$ (for $i\neq j$) occupy incompatible semantic spaces.
@@ 3.2 Stage 1: Training-Free Functional Alignment
Our first stage assembles a MoE model from dense models entirely without training. This process involves two key steps: fusing the shared architectural backbone and functionally aligning the FFN expert layers.
@@
1. Activation Collection: We construct a small and diverse calibration set $D_{\mathrm{cal}}$ by sampling equally from each continue pre-training dataset $D_{1},D_{2},\ldots,D_{N}$} used in dense model training. Each instruction in $D_{\mathrm{cal}}$ is passed through both the anchor model $M_{1}$ and a target model $M_{i}$ to extract their FFN activation matrices at layer $l$, denoted as $A_{1}^{(l)}$ and $A_{i}^{(l)}$, respectively. We sample activations using a tokenizer with a merged vocabulary, extract the post-nonlinearity activation values, and perform normalization.
@@
3. Weight Remapping: We apply the computed permutation $P_{i}^{(l)}$ to the FFN weights (Nasery et al. 2025). This involves permuting the output dimension of the first linear layer’s weights $W_{up}$, $W_{gate}$, and the input dimension of the second linear layer’s $W_{down}$ to maintain harmonization:
$W^{\prime}_{up,i}=W_{up,i}^{(l)}P_{i}^{(l)},\quad W^{\prime}_{gate,i}=W_{gate,i}^{(l)}P_{i}^{(l)}$ (3)
$W^{\prime}_{down,i}=(P_{i}^{(l)})^{T}W_{down,i}^{(l)}$ (4)
@@
We analyzed the computational cost to verify the scalability. The time complexity of this solution is polynomial. The analysis can be found in the Appendix.
3.3 Stage 2: Activating the MoE via Post-training
In Stage 2, we integrate the constructed shared backbone with the expert models and introduce a randomly initialized router, implemented as a simple linear layer. The calibration dataset $D_{\mathrm{cal}}$, introduced in Stage 1, is extended, enabling the MoE to learn how to coordinate the experts. The expert layers, shared backbone, and the router are trained on the extended calibration dataset $D_{cal}$.
We adopt a top-2 routing mechanism and train the MoE model using a composite loss function:
@@ 3.3 Stage 2: Activating the MoE via Post-training
where $\mathcal{L}_{lm}$ is the standard causal language modeling objective, and $\mathcal{L}_{bal}$ is a load-balancing regularization term introduced in Switch Transformer (Fedus, Zoph, and Shazeer 2022). The coefficient $\lambda$ is fixed at 0.01 in all experiments.
Notably, the router is not provided with explicit supervision indicating which expert to select for each token. It learns expert assignment implicitly through continual pre-training on $D_{\mathrm{cal}}$.
@@ 4.1.1 Build Disparately Initialized Dense Models
MMLU GSM8K BBH HumanEval TruthfulQA Avg.* MedCQA
@@
Table 1: Performance comparison of dense (Qwen2/2.5 0.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model. “Avg*” stands for the average scores of ID datasets.
@@
Table 2: Performance comparison of dense (Qwen2/2.5 1.5B) and MoE models (upcycled from dense models) on in-distribution (ID) and out-of-distribution (OOD) data. The best and second-best results are bolded and underlined. “ANC” stands for anchor model. “Avg*” stands for the average scores of ID datasets.
@@ 4.1.2 Evaluation Datasets and Metrics
We evaluate our method on five in-distribution datasets: MMLU (General)  (Hendrycks et al. 2021), GSM8K (Math) (Cobbe et al. 2021), BBH (General, Reasoning)  (Suzgun et al. 2022), HumanEval (Code) (Chen 2021), and TruthfulQA (Science) (Lin, Hilton, and Evans 2022). We also include an out-of-distribution benchmark, MedCQA (Medicine) (Shoham and Rappoport 2024).
@@ 4.2 Implementation Details
Our experiments were conducted on 24 NVIDIA V100 GPUs. The dense model was implemented using LLaMA-Factory (Zheng et al. 2024), and trained for 2 epochs.
In Stage 1: Alignment, we randomly sampled from each training dataset to construct the calibration set $D_{\mathrm{cal}}$, resulting in a total of 10.4M tokens. The training data is sampled from SlimPajama [^1] https://huggingface.co/datasets/cerebras/SlimPajama-627B (General, Code), Finemath [^2] https://huggingface.co/datasets/HuggingFaceTB/finemath, and Scientific Papers [^3] https://huggingface.co/datasets/scientific˙papers. To ensure consistent alignment and integration across expert models, we fixed a general-purpose expert model $M_{1}$ as the anchor model. To ensure that the knowledge of each expert is equally integrated and to avoid introducing any bias, the weights for merging the 4 dense models are set to 0.25.
In Stage 2: Post-training, we increased the sampled data size, expanding $D_{\mathrm{cal}}$ to 5B tokens. We set the cutoff length to 2048 and train the model for 6 epochs with a batch size of 2. The learning rate is fixed at 5e-5, and optimization is performed using AdamW (Zhuang et al. 2022) with a maximum gradient norm of 1.0.
@@ 4.4 Analysis of Inter-Expert Functional Specialization
To quantitatively assess the effectiveness of our functional alignment method, we compute the pairwise functional specialization between expert layers using Centered Kernel Alignment (CKA) (Cortes, Mohri, and Rostamizadeh 2012), a widely adopted technique for comparing neural network representations. A lower CKA score indicates greater functional specialization. We evaluate three scenarios: (a) the original, unmerged expert models as a reference baseline; (b) a naive merging approach without alignment; (c) our method employing activation-based functional alignment.
As shown in Figure 3, the naive merging strategy leads to a severe representational collapse. Experts trained independently develop distinct internal coordinate systems; merging them without alignment forces the model to conflate functionally unrelated neurons. This results in inflated inter-expert CKA scores (e.g., 0.65–0.75), signaling that unique functional fingerprints of the experts have been blurred into a redundant subspace. In such a scenario, the router lacks the discriminative basis to perform specialized dispatching.
In contrast, our method explicitly resolves this misalignment by permuting neurons to form a functionally coherent coordinate system prior to merging. Figure 3(c) demonstrates that this alignment dramatically restores inter-expert specialization, with CKA scores returning to near-optimal levels comparable to those of the original experts (Figure 3(a)). This provides empirical evidence that our harmonization framework effectively mitigates parameter space misalignment.
@@ 4.7 Ablation Analysis
The results highlight the critical role of functional alignment: removing it causes catastrophic performance collapse, confirming that resolving parameter space inconsistencies is fundamental to the framework. Aligning only 80% of the neurons did not result in a significant decrease in the model’s performance, demonstrating the robustness of the alignment method. Replacing SLERP with linear averaging leads to substantial degradation, supporting our hypothesis that the geometry of attention weights is essential for preserving functionality. While simplifying the embedding fusion yields only a moderate drop, it nonetheless indicates the value of layer-aware merging in the shared backbone. Notably, using biased calibration data results in the most severe performance decline after removing alignment, underscoring the importance of domain diversity in constructing $D_{\mathrm{cal}}$ for effective inter-expert coordination.